We collected a sample of the prices of new homes. The mean of our sample is $255,000, with a standard deviation of $19,000. Calculate the z-scores of each of the given prices. Determine if the they are usual or unusual. Round up each z-score to two decimals.

(a) $275, 000

(b) $230,000

(c) $210,000

(d) $290,000
Answer #1

Given u=255000, 6=19000 z score = x-l @ ascore = 275000 - 255000 19000 = 1.05 6 2 score = 230000 - 255000 19ooo = -1.32 z sco